# Fast training and sampling of Restricted Boltzmann Machines

### v2

@@ 3 The Restricted Boltzmann Machine
where $\left\langle{\cdot}\right\rangle_{\mathcal{D}}$ denotes the average with respect to the entries in the dataset, and $\left\langle{\cdot}\right\rangle_{\mathrm{RBM}}$ with respect to $p(\bm{v},\bm{h})$. Since $Z$ is intractable, the model averages in the gradient are typically estimated using $N_{\mathrm{s}}$ independent MCMC processes, and observable averages $\left\langle{o(\bm{v},\bm{h})}\right\rangle_{\mathrm{RBM}}$ are replaced by $\sum_{r=1}^{R}o(\bm{v}^{(r)},\bm{h}^{(r)})/R$, with $(\bm{v}^{(r)},\bm{h}^{(r)})$ being the last configurations reached with each of the $r=1,\ldots,R$ parallel chains. To obtain reliable estimates, it should be ensured that each of the Markov chains mix well before each parameter update. However, ensuring equilibrium at each update is impractical, slow and tedious. The common use of non-convergent MCMC processes is the cause of most difficulties and weird dynamical behaviors encountered in training RBMs as discussed in Decelle et al. (2021). In order to minimize out-of-equilibrium effects, it is often useful to keep $R$ permanent (or persistent) chains, which means that the last configurations reached with the MCMC process used to estimate the gradient at a given parameter update $t$, $\bm{P}_{t}\equiv\{(\bm{v}^{(r)}_{t},\bm{h}^{(r)}_{t})\}_{r=1}^{R}$, are used to initialize the chains of the subsequent update $t+1$. This algorithm is typically referred as PCD. In this scheme, the process of training can be mimicked to a slow cooling process, only that instead of varying a single parameter, e.g. the temperature, a whole set of parameters $\bm{\Theta}_{t}=(\bm{w}_{t},\bm{\theta}_{t},\bm{\eta}_{t})$ are updated at every step to $\bm{\Theta}_{t+1}=\bm{\Theta}_{t}+\gamma\bm{\nabla}\mathcal{L}_{t}$ with $\bm{\nabla}\mathcal{L}_{t}$ being the gradient in equation 2 estimated using the configurations in $\bm{P}_{t}$, and $\gamma$ being the learning rate.
Typical MCMC mixing times in RBMs are short at the beginning of training but increase as it progresses (Decelle et al., 2021). Recent analytical and numerical studies show that the mixing times exhibit sharp, localized peaks when the model encodes a new mode during training (Bachtis et al., 2024). This phenomenon, known in physics as critical slowing down, is related to the system going through critical phase transitions. For RBMs, the first phase transitions during maximum likelihood training are associated with the encoding of the principal components of the dataset into the singular vectors of the weight matrix $\bm{W}$ (Decelle et al., 2017; 2018; Decelle & Furtlehner, 2021b; Bachtis et al., 2024). Since these directions can be determined before training by performing PCA on the data, one can pre-train an RBM to encode them accurately. This avoids crossing the initial phase transitions encountered during the standard training and the associated degradation of the gradient calculation.
4 The low-rank RBM pretrained
In Decelle & Furtlehner (2021a), it was shown that it is possible to train exactly an RBM containing a reduced number of modes in the weight matrix $\bm{W}$ by exploiting a mapping between the RBM and a Restricted Coulomb Machine and solving a convex optimization problem, see the SI A.2. In other words, it is possible to train an RBM whose coupling matrix has this simplified form
$\bm{W}=\sum_{\alpha=1}^{d}w_{\alpha}\bar{\bm{u}}_{\alpha}\bm{u}_{\alpha}^{\top},\qquad\text{with}\qquad(\bm{u}_{\alpha},\bar{\bm{u}}_{\alpha})\in{\mathbb{R}}^{N_{v}}\times{\mathbb{R}}^{N_{h}},$ (3)
where $\{w_{\alpha}\}$ are the singular values of the coupling matrix and the right singular vectors $\{\bm{u}_{\alpha}\}_{\alpha=1}^{d}$ correspond exactly to the first $d$ principal directions of the data set. Under this assumption, $p(\bm{v})$ a function of only $d$ order parameters, given by the magnetizations along each of the $\boldsymbol{u}_{\alpha}$ components, i.e. $m_{\alpha}(\bm{v})=\bm{u}_{\alpha}\cdot\bm{v}/\sqrt{{N_{\mathrm{v}}}}$. In particular,
$H(\bm{v})=-\sum_{\alpha=0}^{d}\theta_{\alpha}m_{\alpha}-\sum_{a}\log\cosh\left(\sqrt{N_{\mathrm{v}}}\bar{u}_{a}\sum_{\alpha=1}^{d}w_{\alpha}m_{\alpha}+\eta_{a}\right)=\mathcal{H}(\bm{m}(\bm{v})),$ (4)
where $\bm{m}=\left(m_{1},\ldots,m_{\alpha}\right)$ and $\theta_{\alpha}$ is the projection of the visible bias on the extended $\bm{u}$ matrix. In order to initialize the visible bias as suggested in Montavon & Müller (2012), a direction $\bm{u}_{0}$ is used for the magnetization $m_{0}$ that is only present in the bias term. This is an improvement of this work with respect to Decelle & Furtlehner (2021a), as it allows us to gain almost for free one additional direction.
The optimal parameters can be determined by solving a convex optimization problem, detailed in SI A.2. Once the model is trained, we obtain a probability distribution $p(\bm{m})$ in a much lower-dimensional space than the original $p(\bm{v})$, which can be easily sampled globally using the inverse transform sampling method. However, since this approach requires discretizing the $\bm{m}$-space for both training and generation, it becomes impractical for intrinsic space dimensions $d>4$ due to memory constraints. These low-rank RBMs are trained to replicate the statistics of the dataset projected onto its first $d$ principal components. Despite their simplicity, the low-rank models can still generate an approximate version of the dataset, as shown in Supplemental Figure 15 for the five datasets previously presented.
In the initial stage of the standard RBM learning process, the model encodes the strongest PCA components of the data through multiple critical transitions (Decelle et al., 2017; 2018; Bachtis et al., 2024). Pre-training with the low-rank construction allows us to bypass these transitions and avoid out-of-equilibrium effects caused by critical slowing down associated to these transitions, which becomes prohibitive in the case of very clustered datasets. Once the main directions are incorporated, training can efficiently continue with standard algorithms like PCD, as the mixing times of pre-trained machines tend to be much shorter that at the transitions.
Figure 2: We compare the equilibrium samples generated by RBMs trained on the MNIST01, Mickey, HGD, Ising2D and CelebA datasets using two different training schemes: PCD, and PCD initialized on low-rank RBMs. To assess the fitting of the modes, we show a scatter plot of the projections of the samples in the first two principal directions of each dataset, together with the histograms along each direction. Black dots and histograms refer to the original data distribution, while the red color is used to represent the generated data. In D we show the evolution of the log-likelihood in time for the two training methods computed with the online AIS trajectory. In continuous lines we show the train and in dashed, the test. Results are obtained by averaging with respect to 10 trainings initialized at the same low-rank RBM, and the standard deviation between all is shown in a shade around the line, which is hardly seen due to the small variation from run to run.
Overcoming these transitions has dramatic implications for the quality of the models we can train and how accurately they reproduce the statistics of the data, especially for short training times. In Fig. 2, we show for 5 datasets the equilibrium samples drawn from 2 RBMs trained with identical number of samples, minibatch size, $k=100$ Gibbs steps, and learning rate $\gamma=0.01$, but different training strategies. In particular, we consider an RBMs trained from scratch with the standard PCD (Tieleman, 2008), and a final machine trained with PCD and pre-trained with a low-rank RBM. We also considered machines trained with the recently proposed Jarzynski reweighing method (Carbone et al., 2023) (see SI E for our specific implementation in the RBM), but we do not include the analysis here because in most of the datasets the training method eventually becomes very unstable reaching very poor results. In all datasets, the quality of the generated samples and the test log-likelihood are significantly better when pre-training is used. We see that the pre-training+PCD training perfectly balances the different modes (here we show the projections on the first 2 PCA directions, but the agreement extends further to several other principal directions). We can also compare the log-likelihood of all 5 datasets and find that the pre-trained RBM achieves higher values. In all cases, the log-likelihood is measured using the annealed importance sampling trajectory (Tr-AIS) method, and the samples are drawn using the trajectory PT (PTT) algorithm. Both algorithms will be described in the next section 5.
Already from the scatter plots, we see that the pre-training is key in obtaining models where all modes are properly balanced in clustered datasets, but it’s also beneficial in the maximum test log-likelihood we can achieve. We observe that the pretraining is not particularly useful when the dataset is not much clustered, as it is the case of CelebA and the entire MNIST dataset as shown in Fig. 16 of the SI I. This result can be easily understood as the encoding of the first modes of the data do not involve splitting the distribution into two isolated modes, something that happens in the case of clustered data.
5 Trajectory annealing
@@
The evolution of parameters during training can be seen as a smoother annealing process compared to the standard temperature annealing used in algorithms like annealing importance sampling (AIS) for log-likelihood estimation or the PT algorithm for sampling. This is because it avoids the first-order transitions that typically occur with temperature in clustered datasets. However, these methods can still be applied to trajectory annealing by using the parameters from different stages of training, rather than scaling the model parameters by a common factor $\beta_{i}$, as is done in temperature-based protocols.
@@ 5.1 Trajectory Annealing importance sampling
The standard method for computing the log-likelihood (LL) in the literature is Annealed Importance Sampling (AIS) (Neal, 2001). For a detailed review of its application in RBMs, see Krause et al. (2020). In this approach, the partition function of a model with parameters $\bm{\Theta}=\{\bm{w},\bm{\theta},\bm{\eta}\}$ is estimated by gradually increasing the inverse temperature $\beta$ from 0 to 1 over $N_{\beta}$ steps. At $\beta=0$ (where $p_{0}=p_{\mathrm{ref}}$), the partition function $Z_{0}$ is tractable, enabling the estimation of $Z_{i}$ at neighboring $\beta_{i}$ values through path sampling methods. The reference configuration $p_{\mathrm{ref}}$ is commonly taken uniform, but improvements have been reported in RBM when fixing it to the independent site distribution that describes the empirical center of the dataset. For a detailed description of the algorithm, see SI D.
@@
To evaluate the reliability of our new approach, we trained several RBMs using datasets with a maximum of 20 hidden nodes, allowing for the exact computation of the partition function through direct enumeration. We then calculated the difference between the true partition function and the estimations obtained from various methods. The results for the HGD dataset are presented in Fig. 3, while similar qualitative results for the other datasets can be found in the SI, see Figs. 11, 12, 13 and 14. In Fig. 3, we display the estimation errors in the LL using three different methods: standard temperature Annealed Importance Sampling (AIS) (panel A), AIS with a reference distribution (panel B), and Tr-AIS (panel C). For the temperature AIS estimates, we show results for different values of $N_{\beta}$ in distinct colors. In panel C, we compare the online estimation in black with the offline Tr-AIS estimation, using the same models analyzed in panels A and B. Additionally, we present the error in LL estimation using a few machines and the trajectory parallel tempering (PTT), which will be detailed in the next section. Overall, the online Tr-AIS estimation consistently outperforms the other methods, and only the PTT estimation yields similar results in offline analyses.
@@
Figure 3: Comparison of the error in (LL) estimation (with respect to the exact value) using different methods: AIS (left), AIS with a reference distribution (middle), and Tr-AIS (ours, right). For Tr-AIS, we present three approaches to computing it: online during training, offline using the models saved during training, and with PTT. The RBM was pretrained and trained with a total of 20 hidden nodes (allowing for an exact computation of the LL) for 10,000 gradient steps. PTT uses a subset of saved models such that the acceptance rate between two consecutive models is 0.25.
5.2 Standard Gibbs sampling vs. Parallel Trajectory Tempering
One major challenge with structured datasets is quantifying the model’s quality, since sampling the equilibrium measure of a well-trained model is often too time-consuming. This affects the reliability of generated samples and indirect measures as log-likelihood’s estimation through Annealing Importance Sampling (AIS) (Krause et al., 2020), making them inaccurate and meaningless.
@@
Figure 4: Comparison of the performance of different MCMC sampling methods on RBMs trained with the pretraining+PCD procedure on the MNIST01 (row-1), the HGD (row-2) and Ising 2D datasets (row-3). In A and B columns, we show the trajectory of two independent Markov chains (red and orange) iterated for $10^{4}$ MCMC steps using either PTT or alternate Gibbs sampling (AGS), projected onto the first two principal components of the dataset. The position of the chains is plotted every 10 MCMC steps. The black contour represents the density profile of the dataset. In column C, we show the average number of jumps between the two regions separated by the dashed grew line in the PCA plots using different MCMC methods: Alternative Gibbs Sampling (AGS), Parallel Tempering (PT) (Hukushima & Nemoto, 1996), without or with reference configuration (PTref) (Krause et al., 2020), the Stacked Tempering (Roussel et al., 2023), and the Trajectory Parallel Tempering (PTT) proposed in this work. The average is calculated over a population of 1000 chains and the shadow around the lines indicates the error of the mean.
@@
To illustrate this problem, let us consider the MNIST01, the HGD and the Ising 2D datasets. MNIST01 and Ising 2D are strongly bimodal in the PCA and the HGD trimodal, as shown in Fig. 1. Let us consider that we want to sample the equilibrium measure of the RBMs trained using low-rank RBM pretraining. In order to draw new samples from these models, one would typically run MCMC processes from random initialization and iterate them until convergence. The mixing time in such case is controlled by the jumping time between clusters. To accurately estimate the relative weight between modes, the MCMC processes must be ergodic, requiring many back-and-forth jumps. However, as shown in Figs. 4–A1-3, Alternate Gibbs sampling (AGS) dynamics are extremely slow, rarely producing jumps even after $10^{4}$ MCMC steps. The red curves in Figs. 4–C1-3 show the mean number of jumps over 1000 independent chains as a function of MCMC steps, indicating that a proper equilibrium generation would require at least $10^{6}-10^{7}$ MCMC steps. In order to speed up sampling, we consider other traditional optimized MCMC methods such as the PT (with or without reference), the Stack Tempering (Roussel et al., 2023) and our Parallel Trajectory Tempering (PTT) that will be explained below. The evolution of the number of jumps between modes observed with each of these algorithms, as a function of the total number of AGS steps performed (i.e. the number of parallel models multiplied by the number of AGS steps performed at each model) is shown in Fig. 4–C1-3. Before discussing our PTT algorithm and the results, let us briefly introduce the PT algorithm.
5.2.1 Parallel Tempering approaches
@@
Parallel tempering (PT) is commonly used to simulate multiple systems in parallel at different inverse temperatures $\beta$ (Hukushima & Nemoto, 1996; Marinari & Parisi, 1992), with occasional temperature swaps between models based on the Metropolis rule. Since this move satisfies detailed balance, the dynamics will converge to the Boltzmann distribution $p_{\beta}(\bm{x})\propto\exp(-\beta\mathcal{H}(\bm{x}))$ at each $\beta$. However, a more general approach involves simulating multiple systems, each with its own set of parameters. We can define a system as $p_{\bm{\Theta}^{t}}(\bm{x})$, where $\bm{\Theta}^{t}$ denotes the parameters of the model indexed by $t$. This includes the classic PT case, where temperature scaling is applied, $\bm{\Theta}^{t}=\beta^{(t)}\bm{\Theta}$, with $\bm{\Theta}$ being the parameters of the model to be sampled.
@@
In this work, we introduce a new algorithm that exchanges parameters of models saved at different stages of learning, using the learning trajectory to define the parameter set $\bm{\Theta}^{t}$, where $t$ refers to a specific point in the training process. We call this method Parallel Trajectory Tempering (PTT). The rule for swapping configurations $\bm{x}=(\bm{v},\bm{h})$ between neighboring machines indexed by $t$ and $t-1$ is given by
@@
$p_{\mathrm{acc}}(\bm{x}^{t}\leftrightarrow\bm{x}^{t-1})=\min\left(1,\exp\left(\Delta\mathcal{H}^{t}(\bm{x}^{t})-\Delta\mathcal{H}^{t}(\bm{x}^{t-1})\right)\right),$ (5)

### v3

@@ 3 The Restricted Boltzmann Machine
Since $Z$ is intractable, gradient estimates typically rely on $N_{\mathrm{s}}$ independent MCMC processes, replacing observable averages $\left\langle{o(\bm{v},\bm{h})}\right\rangle_{\mathrm{RBM}}$ with $\sum_{r=1}^{R}o(\bm{v}^{(r)},\bm{h}^{(r)})/R$, where $(\bm{v}^{(r)},\bm{h}^{(r)})$ are the final states of $R$ parallel chains. Reliable estimates require proper mixing before each parameter update, but enforcing equilibrium at every step is impractical and slow. The widespread use of non-convergent MCMC processes underlies most training difficulties and anomalous dynamics in RBMs, as discussed in (decelle2021equilibrium). In order to minimize out-of-equilibrium effects, it is often useful to keep $R$ permanent (or persistent) chains, which means that the last configurations reached with the MCMC process used to estimate the gradient at a given parameter update $t$, $\bm{P}_{t}\equiv\{(\bm{v}^{(r)}_{t},\bm{h}^{(r)}_{t})\}_{r=1}^{R}$, are used to initialize the chains of the subsequent update $t+1$. This algorithm is typically referred to as PCD. In this scheme, the process of training can be mimicked as a slow cooling process, only that instead of varying a single parameter, e.g. the temperature, a whole set of parameters $\bm{\Theta}_{t}=(\bm{w}_{t},\bm{\theta}_{t},\bm{\eta}_{t})$ is updated at every step to $\bm{\Theta}_{t+1}=\bm{\Theta}_{t}+\gamma\bm{\nabla}\mathcal{L}_{t}$ with $\bm{\nabla}\mathcal{L}_{t}$ being the gradient in equation 2 estimated using the configurations in $\bm{P}_{t}$, and $\gamma$ being the learning rate.
Typical MCMC mixing times in RBMs are initially short but increase as training progresses (decelle2021equilibrium). Studies reveal sharp mixing time peaks when the model encodes new modes (bachtis2024cascade), a phenomenon known as critical slowing down hohenberg1977theory, associated with critical phase transitions. In RBMs, these transitions are related with the alignment of the singular vectors of $\bm{W}$ with the dataset’s principal components (Decelle_2017; decelle2018thermodynamics; decelle2021restricted; bachtis2024cascade). However, these dominant modes can be identified via PCA before training and efficiently encoded through convex optimization, circumventing standard maximum likelihood training. This approach prevents gradient degradation due to diverging mixing times at the first second-order transitions. The advantages of using the low-rank construction from decelle2021exact as a pretraining method will be discussed in Section 5.
4 Trajectory annealing paradigm
@@
One major challenge with structured datasets is quantifying the model’s quality, since sampling the equilibrium measure of a well-trained model is often too time-consuming. This affects the reliability of generated samples and indirect measures such as log-likelihood (LL) estimation through Annealing Importance Sampling (AIS) (krause_algorithms_2020), making them inaccurate and meaningless. The evolution of parameters during training can be seen as a smoother annealing process compared to the standard temperature annealing used in algorithms like annealing importance sampling (AIS) for LL estimation or in sampling algorithms like simulated annealing or PT. This is because it avoids the first-order transitions that typically occur with temperature in clustered datasets, see SI H for more detailed explanations of this phenomenon in clustered datasets. However, these methods can still be applied to trajectory annealing using the parameters of different stages of training, rather than scaling the model parameters by a common factor $\beta_{i}$, as is done in temperature-based protocols.
@@ 4.1 Trajectory Annealing importance sampling
The standard method for log-LL estimation in the literature is AIS (neal2001annealed); see (krause_algorithms_2020) for a review on its application to RBMs. A detailed algorithm description is provided in SI D. AIS estimates the partition function of a model with parameters $\bm{\Theta}=\{\bm{W},\bm{\theta},\bm{\eta}\}$ by gradually increasing the inverse temperature $\beta$ from 0 to 1 over $N_{\beta}$ steps. At $\beta=0$, where $p_{0}=p_{\mathrm{ref}}$, the partition function $Z_{0}$ is tractable, allowing estimation of $Z_{i}$ at neighboring $\beta_{i}$ values via path sampling. The reference distribution $p_{\mathrm{ref}}$ is typically uniform, but improved performance has been reported when using the independent site distribution centered on the dataset, see SI D for details.
@@
To evaluate the reliability of our new approach, we trained several RBMs using datasets with a maximum of 20 hidden nodes, allowing for the exact computation of the partition function through direct enumeration. We then calculated the difference between the true partition function and the estimations obtained from various methods. The results for the HGD dataset are presented in Fig. 2, while similar qualitative results for the other datasets can be found in the SI, see Figs. 15, 16, 17 and 18. In Fig. 2, we display the estimation errors in the LL using three different methods: standard temperature AIS (panel A), AIS with non-flat reference distribution (panel B), and Tr-AIS (panel C). For the temperature AIS estimates, we show results for different values of $N_{\beta}$ in distinct colors. In panel C, we compare the online estimation in black with the offline Tr-AIS estimation, using the same models analyzed in panels A and B. Additionally, we present the error in LL estimation using just a few machines saved during the training and the trajectory parallel tempering (PTT), which will be detailed in the next section. Overall, the online Tr-AIS estimation consistently outperforms the other methods, and only the PTT estimation yields similar results in offline analyses.
@@
Figure 2: Comparison of LL estimation error (relative to the exact value) across different methods: AIS (A), AIS with a reference distribution fixed to the independent site distribution matching the dataset’s empirical center (B), and Tr-AIS (ours, C). For Tr-AIS, we evaluate three approaches: online during training, offline using saved models, and with PTT. The RBM, pretrained and trained with 20 hidden nodes (allowing exact $L$ computation), for $10^{4}$ gradient steps on the HGD dataset. PTT selects a subset of saved models ensuring a 0.25 acceptance rate between consecutive models. Lines show the mean $L$ over 10 independent runs, with shaded areas representing one standard deviation.
4.2 Exploiting the training trajectory for sampling
The trajectory annealing paradigm can also be exploited to significantly accelerate the sampling of equilibrium configurations, as we show in Fig. 3. Before presenting our parallel trajectory annealing (PTT) algorithm, we first discuss the limitations of standard methods like Alternating Gibbs Sampling (AGS), and introduce the standard parallel tempering (PT) algorithm. The AGS procedure involves iteratively alternating between two steps: conditioning on all hidden variables given fixed visible variables, and then conditioning on all visible variables given fixed hidden variables.
@@ 4.2 Exploiting the training trajectory for sampling
Figure 3: Comparison of the performance of different MCMC sampling methods on RBMs trained with the pretraining+PCD procedure on the MNIST01 (row-1), the HGD (row-2) and Ising 2D datasets (row-3). In A and B columns, we show the trajectory of two independent Markov chains (red and orange) iterated for $10^{4}$ MCMC steps using either PTT or AGS, projected onto the first two principal components of the dataset. The position of the chains is plotted every 10 MCMC steps. The black contour represents the density profile of the dataset. In column C, we show the averaged number of jumps between the two regions separated by the dashed grey line in the PCA plots using different $MC$ methods: AGS, Parallel Tempering (PT) (hukushima1996exchange), without or with reference configuration (PTref) (krause_algorithms_2020), the Stacked Tempering (roussel2023accelerated), and the Trajectory Parallel Tempering (PTT) proposed in this work. The average is calculated over a population of 1000 chains and the shadow around the lines indicates the error of the mean.
@@
Problems with the standard AGS method– Let us consider the MNIST01, HGD, and Ising 2D datasets. MNIST01 and Ising 2D exhibit strong bimodality in their PCA representations, while HGD is trimodal, as shown in Fig. 1. Suppose that we have trained an RBM on each of these datasets, for example using the low-rank RBM pretraining (see Section 5), and we aim to sample the equilibrium measure of these models. Typically, this involves running MCMC processes from random initializations and iterating them until convergence. The mixing time in such cases is determined by the jumping time between clusters. Accurate estimation of the relative weights between modes requires the MCMC processes to be ergodic, necessitating frequent back-and-forth jumps. However, as shown in the chains in Figs. 3–B1-3, AGS dynamics are exceedingly slow, rarely producing jumps even after $10^{4}$ MCMC steps. The red curves in Figs. 3–C1-3 illustrate the mean number of jumps over 1000 independent chains as a function of MCMC steps, indicating that achieving proper equilibrium generation would require at least $10^{6}$-$10^{7}$ MCMC steps. To accelerate sampling, we therefore explore optimized Monte Carlo methods inspired by Parallel Tempering.
@@
Trajectory Parallel Tempering– Parallel tempering (PT) is used to simulate multiple systems in parallel at different inverse temperatures $\beta$ (hukushima1996exchange; marinari1992simulated), with occasional temperature swaps between models based on the Metropolis rule. Since this move satisfies detailed balance, the dynamics will converge to the Boltzmann distribution $p_{\beta}(\bm{x})\propto\exp(-\beta\mathcal{H}(\bm{x}))$ at each $\beta$. However, a more general approach involves simulating multiple systems with distincts set of parameters. For a set of parameters $\bm{\Theta}^{t}$ we associate the system $p_{\bm{\Theta}^{t}}(\bm{x})$. This includes the classic PT case, with $\bm{\Theta}_{\mathrm{PT}}^{t}=\beta{(t)}\bm{\Theta}$ where the temperature ladder is typically defined as $\beta(t)=t/N_{T}$, where $t=0,\ldots,N_{T}-1$, and $N_{T}$ represents the number of systems simulated in parallel. Alternatively, recent work (roussel2023accelerated) introduced a variant of this algorithm called Stacked Tempering, which uses a series of nested RBMs, where the hidden configurations of one machine are exchanged with the visible configurations of the next in the stack, analogous to the mechanism in deep belief networks.
@@
We introduce a novel variant where model parameters from different training phases are exchanged. Using the learning trajectory, we define the parameter set $\bm{\Theta}^{t}$, with $t$ indexing models at specific training steps. We call this method Parallel Trajectory Tempering (PTT). Configuration exchanges $\bm{x}\!=\!(\bm{v},\bm{h})$ between neighboring models at $t$ and t1t\!-\!1 follow the rule:
@@
$p_{\mathrm{acc}}(\bm{x}^{t}\!\leftrightarrow\!\bm{x}^{t-1})\!=\!\min\left[1,\exp\left(\Delta\mathcal{H}^{t}(\bm{x}^{t})\!-\!\Delta\mathcal{H}^{t}(\bm{x}^{t-1})\right)\right]\>\>\text{with}\>\>\Delta\mathcal{H}_{t}(\bm{x})\!\equiv\!\mathcal{H}_{t}(\bm{x})-\mathcal{H}_{t-1}(\bm{x}),$ (3)